# ReAct Meets ActRe: Autonomous Annotations of Agent Trajectories for Contrastive Self-Training

### v1

@@ 5 Conclusion
In this work, we propose AT, a framework that enables the autonomous annotations of agent trajectories in the style of ReAct for contrastive self-training. The key factor in the trajectory annotation process is the ActRe prompting agent, which produces the textual rationales given arbitrary external actions. Together with ActRe and environmental feedback, the ReAct-style agent autonomously synthesizes trajectories for self-training. In the contrastive self-training process, we leverage the policy gradient methods with binarized rewards to boost the task success rate. Extensive experiments on AlfWorld and WebShop have demonstrated the superiority of AT over multiple strong baselines.

### v2

@@ 5 Conclusion
In this work, we propose A3T, a framework that enables the autonomous annotation of agent trajectories in the style of ReAct for contrastive self-training. The key factor in the trajectory annotation process is the ActRe prompting agent, which produces the textual rationales given arbitrary external actions. Together with ActRe and environmental feedback, the ReAct-style agent autonomously synthesizes trajectories for self-training. In the contrastive self-training process, we leverage the policy gradient methods with binarized rewards to boost the task success rate. Extensive experiments on AlfWorld and WebShop have demonstrated the superiority of A3T over multiple strong baselines.